The correct order of bond length for CO, CO2, CO3 2-(carbonate) is:


why?

Grade:11

5 Answers

SAGAR SINGH - IIT DELHI
12 years ago

When we consider a bond between two atoms, the type of bond gives a relative measure of the bond length.

Triple Bond Strongest, 3 electron pairs shared, shortest bond
Double Bond Middle, 2 electron pairs shared
Single Bond Weakest, 1 electron pair shared, longest bond


We need to determine the bond types for each molecule or ion:

CO, When drawing the lewis structure here, we end up with a triple bond between the carbon and the oxygen (and a -1 charge on the C and a +1 charge on the O)

CO2 This lewis structure gives one double bond between each oxygen and the carbon.

[CO3]2- This one is a little trickier, but when we draw the lewis structure for this ion, we get a double bond between one oxygen and the carbon, with single bonds between the other two oxygen atoms and the carbon. Because each of the oxygens want a double bond to the carbon, they all fight over it (giving resonance structures), meaning that in the end, you get a bond something between a double and a single bond for each Carbon-Oxygen bond. Therefore, this bond will be longer than a single bond, but shorter than a double bond:

So, this give the following increasing order of bond length:

[CO3]2-, CO2, CO

Darryl
6 years ago
Actually for resonance structures, it’s intermediate between a single bond and double bond
 
[CO3]2- 
 
The bond will be shorter than a single bond, but longer than a double bond. 
Sachin Jain
6 years ago
Bond length is inversely proportional to bond order. Bond order of c-o in CO, CO2 and Co3-2 is 3,2,1.33 respectively.
Shivraj
5 years ago
Bond length is inversly proportional to bond order 
Bond order of CO calculated by molecular orbital theory =3
B.O. of CO2 =no. Of bonds in all possible sides / no. Of resoniting structures =4/2=2 
Similarly B.O. of CO3 -2 =4/3 1.33
Therefor bond length of CO
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

CO<CO2<CO3^2- is the increasing bond length order because
CO has triple bond so bond length minimum and CO2 has double bond so its bond length maximum then CO and CO32- has partial single bond characters due to resonance so it has maximum bond lenght

Thanks and Regards